# HOG-Diff: Higher-Order Guided Diffusion for Graph Generation

### v1

@@ D.2 Training and Sampling Proceudre
Input: denoising network $\bm{s}_{\bm{\theta}}^{(k)}$, authentic graph data $\bm{G}_{0}=(\bm{X}_{0},\bm{A}_{0})$, cell complex list $\{\mathcal{S}^{(0)},\cdots,\mathcal{S}^{(K)}\}$, - training epochs $M_{k}$.
@@
4:    $\bm{U}_{0}\leftarrow\operatorname{EigenVectors}(\bm{D}_{0}-\bm{A}_{0})$
5:    $\bm{\Lambda}_{\tau_{k}}\leftarrow\operatorname{EigenDecompostion}(\bm{D}_{\tau_{k}}-\bm{A}_{\tau_{k}})$
6:    $\bm{\Lambda}_{\tau_{k-1}}\leftarrow\operatorname{EigenDecompostion}(\bm{D}_{\tau_{k-1}}-\bm{A}_{\tau_{k-1}})$
7:    Sample $t\sim\operatorname{Unif}([0,\tau_{k}-\tau_{k-1}])$
@@
10:    $\bm{L}_{t}^{(k)}\leftarrow\bm{U}_{0}\bm{\Lambda}_{t}^{(k)}\bm{U}_{0}^{\top}$
11:    $\bm{A}_{t}^{(k)}\leftarrow\bm{D}_{t}^{(k)}-\bm{L}_{t}^{(k)}$
12:    $\ell^{(k)}(\bm{\theta})\leftarrow c_{1}\|\bm{s}^{(k)}_{\bm{\theta},\bm{X}}-\nabla_{\bm{X}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})\|^{2}+c_{2}||\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}-\nabla_{\bm{\Lambda}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})||^{2}$
13:    $\bm{\theta}\leftarrow\operatorname{optimizer}(\ell^{(k)}(\bm{\theta}))$
14:  end for
15:  Return: $\bm{s}^{(k)}_{\bm{\theta}}$
The pseudo-code of sampling with HOG-Diff is described in Alg. 2. The reverse diffusion processes are divided into $K$ hierarchical time windows, denoted as $\{[\tau_{k-1},\tau_{k}]\}_{k=1}^{K}$, where $0=\tau_{0}<\cdots<\tau_{k-1}<\tau_{k}<\cdots<\tau_{K}=T$. We first initialize the sampling process by drawing samples for $\widehat{\bm{X}}_{\tau_{K}}$ and $\widehat{\bm{\Lambda}}_{\tau_{K}}$ from a standard Gaussian distribution, and $\widehat{\bm{U}}_{0}$ is sampled uniformly from the eigenvector matrices of the Laplacian matrix in the training dataset. The reverse-time process starts at $\tau_{K}$ and iteratively updates $\widehat{\bm{X}}_{t}$ and $\widehat{\bm{\Lambda}}_{t}$ by solving the reverse-time SDEs with the denoising network $\bm{s_{\theta}}^{(k)}$. Subsequently, we reconstruct the Laplacian matrix $\widehat{\bm{L}}_{t}$ using the fixed eigenvector matrix $\widehat{\bm{U}}_{0}$ and the updated eigenvalues $\widehat{\bm{\Lambda}}_{t}$. Endpoint of one generation step serves as the starting point for the next process. Finally, after iterating through all diffusion segments, the algorithm returns the final feature matrix $\widehat{\bm{X}}_{0}$ and adjacency matrix $\widehat{\bm{A}}_{0}$, thereby completing the graph generation process.
@@
Input: Trained denoising network $\bm{s}_{\theta}^{(k)}$, diffusion time split $\{\tau_{0},\cdots,\tau_{K}\}$, number of sampling steps $M_{k}$
1:  $t\leftarrow\tau_{K}$
2:  $\widehat{\bm{X}}_{\tau_{K}}\sim\mathcal{N}(\bm{0},\bm{I})$ and $\widehat{\bm{\Lambda}}_{\tau_{K}}\sim\mathcal{N}(\bm{0},\bm{I})$
3:  $\widehat{\bm{U}}_{0}\sim\operatorname{Unif}\left(\{\bm{U}_{0}\triangleq\operatorname{EigenVectors}(\bm{L}_{0})\}\right)$
4:  $\widehat{\bm{G}}_{\tau_{K}}\leftarrow(\widehat{\bm{X}}_{\tau_{K}},\widehat{\bm{\Lambda}}_{\tau_{K}},\widehat{\bm{D}}_{\tau_{K}}-\widehat{\bm{U}}_{0}\widehat{\bm{\Lambda}}_{\tau_{K}}\widehat{\bm{U}}_{0}^{\top})$
5:  for $k=K$ to $1$ do
6:    for $m=M_{k}-1$ to 00 do
@@
Appendix E Complexity Analysis
When the targeted graph is not in the desired higher-order forms, one should also consider the one-time preprocessing procedure for graph filtering.
Cell filtering can be dramatically accelerated because it avoids explicitly finding all cells and only determines whether nodes and edges belong to a cell. Specifically, the $2$-cell filter requires only checking whether each edge belongs to some cycle.
One method to achieve the $2$-cell filter is using a depth-first search (DFS). Starting from the adjacency matrix, we temporarily remove the edge $(i,j)$ and initiate a DFS from node $i$, keeping track of the path length. If the target node $j$ is visited within a path length of $l$, the edge $(i,j)$ is marked as belonging to a $2$-cell of length at most $l$. In sparse graphs with $n$ nodes and $m$ edges, the time complexity of a single DFS is $\mathcal{O}(m+n)$. With the path length limited to $l$, the DFS may traverse up to $l$ layers of recursion in the worst case. Therefore, the complexity of a single DFS is $\mathcal{O}(\min(m+n,l\cdot k_{max}))$, where $k_{max}$ is the maximum degree of the graph. For all $m$ edges, the total complexity is $\mathcal{O}\left(m\cdot\min(m+n,l\cdot k_{max})\right)$.
@@ Appendix E Complexity Analysis
Appendix F Experimental Setup
F.1 Computing Resources
@@
In this work, all experiments are conducted using PyTorch on a single NVIDIA L40S GPU with 46 GB memory and AMD EPYC 9374F 32-Core Processor.
F.2 Generic Graph Generation
We follow the experimental and evaluation setting from Jo et al. (2022) with the same train/test split to ensure a fair comparison with baselines. We use node degree and spectral features of the graph Laplacian decomposition as hand-crafted input features.
@@ F.2 Generic Graph Generation
Tab. 3 summarizes the key characteristics of the datasets utilized in this study. The table outlines the type of dataset, the total number of graphs, and the range of graph sizes ($|V|$). Additionally, it also provides the number of distinct node types and edge types for each dataset. Notably, the synthetic datasets (Community-small and Ego-small) contain relatively small graphs, whereas the molecular datasets (QM9 and ZINC250k) exhibit more diversity in graph size and complexity.
@@
Dataset Graph type #Graphs #Nodes Node types Edge types
@@
F.3 Molecule Generation
In experiments, each molecule is preprocessed into a graph comprising adjacency matrix $\bm{A}\in\{0,1,2,3\}^{n\times n}$ and node feature matrix $\bm{X}\in\{0,1\}^{n\times d}$, where $n$ denotes the maximum number of atoms in a molecule of the dataset, and $d$ is the number of possible atom types. The entries of $\bm{A}$ indicate the bond types: 0 for no bound, 1 for the single bond, 2 for the double bond, and 3 for the triple bond. Further, we scale $\bm{A}$ with a constant scale of 3 in order to bound the input of the model in the interval [0, 1], and rescale the final sample of the generation process to recover the bond types. Following the standard procedure (Shi et al., 2020; Luo et al., 2021), all molecules are kekulized by the RDKit library (Landrum et al., 2016) with hydrogen atoms removed. In addition, we make use of the valency correction proposed by Zang & Wang (2020). After generating samples by simulating the reverse diffusion process, the adjacency matrix entries are quantized to discrete values ${0,1,2,3}$ by by applying value clipping. Specifically, values in $(-\infty,0.5)$ are mapped to 0, $[0.5,1.5)$ to 1, $[1.5,2.5)$ to 2, and $[2.5,+\infty)$ to 3, ensuring the bond types align with their respective categories.
To comprehensively assess the quality of the generated molecules across datasets, we evaluate 10,000 generated samples using several key metrics: validity, validity w/o check, Frechet ChemNet Distance (FCD) (Preuer et al., 2018), Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) MMD (Costa & Grave, 2010), uniqueness, and novelty (Jo et al., 2022). FCD quantifies the similarity between generated and test molecules by leveraging the activations of ChemNet’s penultimate layer, accessing the generation quality within the chemical space. In contrast, NSPDK-MMD evaluates the generation quality from the graph topology perspective by computing the MMD between the generated and test sets while considering both node and edge features. Validity is measured as the fraction of valid molecules to all generated molecules after applying post-processing corrections such as valency adjustments or edge resampling, while validity w/o correction, following Jo et al. (2022), computes the fraction of valid molecules before any corrections, providing insight into the intrinsic quality of the generative process. Whether molecules are valid is generally determined by compliance with the valence rules in RDkit (Landrum et al., 2016). Novelty assesses the model’s ability to generalize by calculating the percentage of generated graphs that are not subgraphs of the training set, with two graphs considered identical if isomorphic. Uniqueness quantifies the diversity of generated molecules as the ratio of unique samples to valid samples, removing duplicates that are subgraph-isomorphic, ensuring variety in the output.
Appendix G Visualization Results

### v2

@@ C.4 Training and Sampling Procedure
Input: Score network $\bm{s}_{\bm{\theta}}^{(k)}$, training graph dataset $\mathcal{G}$, training epochs $M_{k}$.
@@
5:  $\bm{U}_{0}\leftarrow\operatorname{EigenVectors}(\bm{D}_{0}-\bm{A}_{0})$
6:  $\bm{\Lambda}_{\tau_{k}}\leftarrow\operatorname{EigenDecomposition}(\bm{D}_{\tau_{k}}-\bm{A}_{\tau_{k}})$
7:  $\bm{\Lambda}_{\tau_{k-1}}\leftarrow\operatorname{EigenDecomposition}(\bm{D}_{\tau_{k-1}}-\bm{A}_{\tau_{k-1}})$
8:  Sample $t\sim\operatorname{Unif}([0,\tau_{k}-\tau_{k-1}])$
@@
11:  $\bm{L}_{t}^{(k)}\leftarrow\bm{U}_{0}\bm{\Lambda}_{t}^{(k)}\bm{U}_{0}^{\top}$
12:  $\bm{A}_{t}^{(k)}\leftarrow\bm{D}_{t}^{(k)}-\bm{L}_{t}^{(k)}$
13:  $\ell^{(k)}(\bm{\theta})\leftarrow c_{1}\|\bm{s}^{(k)}_{\bm{\theta},\bm{X}}-\nabla_{\bm{X}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})\|^{2}+c_{2}||\bm{s}^{(k)}_{\bm{\theta},\bm{\Lambda}}-\nabla_{\bm{\Lambda}}\log p_{t}(\bm{G}_{t}|\bm{G}_{\tau_{k}})||^{2}$
14:  $\bm{\theta}\leftarrow\operatorname{optimizer}(\ell^{(k)}(\bm{\theta}))$
15:end for
16:Return: $\bm{s}^{(k)}_{\bm{\theta}}$
@@
Input: Trained score network $\bm{s}_{\theta}^{(k)}$, diffusion time split $\{\tau_{0},\cdots,\tau_{K}\}$, number of sampling steps $M_{k}$
@@
[…]
5:for $k=K$ to 11 do
6:  for $m=M_{k}-1$ to 0 do
8:   $\widehat{\bm{X}}_{t}\leftarrow\widehat{\bm{X}}_{t}-\left[\theta_{t}\left(1+\frac{2}{e^{2\bar{\theta}_{t:\tau_{k}}}-1}\right)(\widehat{\bm{X}}_{\tau_{k}}-\widehat{\bm{X}}_{t})-g_{k,t}^{2}\bm{S_{X}}\right]\delta t+g_{k,t}\sqrt{\delta t}\bm{w}_{\bm{X}}$, $\bm{w}_{\bm{X}}\sim\mathcal{N}(\bm{0},\bm{I})$
9:   $\widehat{\bm{\Lambda}}_{t}\leftarrow\widehat{\bm{\Lambda}}_{t}-\left[\theta_{t}\left(1+\frac{2}{e^{2\bar{\theta}_{t:\tau_{k}}}-1}\right)(\widehat{\bm{\Lambda}}_{\tau_{k}}-\widehat{\bm{\Lambda}}_{t})-g_{k,t}^{2}\bm{S_{\Lambda}}\right]\delta t+g_{k,t}\sqrt{\delta t}\bm{w}_{\bm{\Lambda}}$, $\bm{w}_{\bm{\Lambda}}\sim\mathcal{N}(\bm{0},\bm{I})$
@@
Appendix D Complexity Analysis
When the targeted graph is not in the desired higher-order forms, one should also consider the one-time preprocessing procedure for cell filtering. Unlike cell lifting, which enumerates all cell structures and can incur substantial computational overhead, cell filtering can be performed much more efficiently. This is because filtering does not require the explicit enumeration of all cells; instead, it only checks whether individual nodes and edges participate in a cell. For instance, the $2$-cell filter requires only checking whether each edge belongs to some cycle.
One method to achieve the $2$-cell filter is to use a depth-first search (DFS) strategy. Starting from the adjacency matrix, we temporarily remove the edge $(i,j)$ and initiate a DFS from node $i$, keeping track of the path length. If the target node $j$ is visited within a path length of $l$, the edge $(i,j)$ is marked as belonging to a $2$-cell of length at most $l$. In sparse graphs with $n$ nodes and $m$ edges, the time complexity of a single DFS is $\mathcal{O}(m+n)$. With the path length limited to $l$, the DFS may traverse up to $l$ layers of recursion in the worst case. Therefore, the complexity of a single DFS is $\mathcal{O}(\min(m+n,l\cdot k_{max}))$, where $k_{max}$ is the maximum degree of the graph. For all $m$ edges, the total complexity is $\mathcal{O}\left(m\cdot\min(m+n,l\cdot k_{max})\right)$.
@@ Appendix D Complexity Analysis
After preprocessing, the training loop is identical to that of standard diffusion-based generative models. The proposed coarse-to-fine strategy splits a long diffusion trajectory into shorter segments, each trained with a smaller smoothness constant, which by Theorem 3 provably leads to faster convergence than classical models in idealized settings. Furthermore, since the sub-processes are independent, they can be trained in parallel for additional efficiency gains. In practice, we observe no measurable slowdown relative to GDSS or DiGress.
For inference, Tab. 4 shows that the extra guidance logic does not slow sampling. HOG-Diff matches DiGress on smaller graphs and achieves up to 12×speedup over both baselines on the protein-scale Enzymes dataset.
In short, the only scaling cost is a one-off preprocessing pass; both training and sampling remain GPU-bound and on par with, or even faster than, existing diffusion models.
Table 4: Sampling time (s).
Method Community-small Enzymes Ego-small
GDSS 41 110 28
DiGress 8 301 13
HOG-Diff 11 26 15
Appendix E Experimental Setup
E.1 Computing Resources
@@
In this work, all experiments are conducted using PyTorch (Paszke, 2019) on a single NVIDIA L40S GPU with 46 GB memory and AMD EPYC 9374F 32-Core Processor.
E.2 Molecule Generation
We evaluate the quality of generated molecules on two well-known molecular datasets: QM9 (Ramakrishnan et al., 2014) and ZINC250k (Irwin et al., 2012), and obtain the intermediate higher-order skeletons using the 2-cell complex filtering. In experiments, each molecule is preprocessed into a graph comprising adjacency matrix $\bm{A}\in\{0,1,2,3\}^{n\times n}$ and node feature matrix $\bm{X}\in\{0,1\}^{n\times d}$, where $n$ denotes the maximum number of atoms in a molecule of the dataset, and $d$ is the number of possible atom types. The entries of $\bm{A}$ indicate the bond types: 0 for no bound, 1 for the single bond, 2 for the double bond, and 3 for the triple bond. Further, we scale $\bm{A}$ with a constant scale of 3 in order to bound the input of the model in the interval [0, 1], and rescale the final sample of the generation process to recover the bond types. Following the standard procedure (Shi et al., 2020; Luo et al., 2021), all molecules are kekulized by the RDKit library (Landrum et al., 2016) with hydrogen atoms removed. In addition, we make use of the valency correction proposed by Zang & Wang (2020). After generating samples by simulating the reverse diffusion process, the adjacency matrix entries are quantized to discrete values 0,1,2,3{0,1,2,3} by by applying value clipping. Specifically, values in $(-\infty,0.5)$ are mapped to 0, $[0.5,1.5)$ to 1, $[1.5,2.5)$ to 2, and $2.5,+\infty)$ to 3, ensuring the bond types align with their respective categories.
To comprehensively assess the quality of the generated molecules across datasets, we evaluate 10,000 generated samples using several key metrics: validity, validity w/o check, Frechet ChemNet Distance (FCD) (Preuer et al., 2018), Neighborhood Subgraph Pairwise Distance Kernel (NSPDK) MMD (Costa & Grave, 2010), uniqueness, and novelty (Jo et al., 2022). FCD quantifies the similarity between generated and test molecules by leveraging the activations of ChemNet’s penultimate layer, assessing the generation quality within the chemical space. In contrast, NSPDK-MMD evaluates the generation quality from the graph topology perspective by computing the MMD between the generated and test sets while considering both node and edge features. Validity is measured as the fraction of valid molecules to all generated molecules after applying post-processing corrections such as valency adjustments or edge resampling, while validity w/o correction, following Jo et al. (2022), computes the fraction of valid molecules before any corrections, providing insight into the intrinsic quality of the generative process. Whether molecules are valid is generally determined by compliance with the valence rules in RDKit (Landrum et al., 2016). Novelty assesses the model’s ability to generalize by calculating the percentage of generated graphs that are not subgraphs of the training set, with two graphs considered identical if isomorphic. Uniqueness quantifies the diversity of generated molecules as the ratio of unique samples to valid samples, removing duplicates that are subgraph-isomorphic, ensuring variety in the output. We report the baseline results taken from Jo et al. (2022) and Kong et al. (2023).
E.3 Generic Graph Generation
To display the topology distribution learning ability, we assess HOG-Diff over four common generic graph datasets: (1) Community-small, containing 100 randomly generated community graphs; (2) Ego-small, comprising 200 small ego graphs derived from the Citeseer network dataset; (3) Enzymes, featuring 587 protein graphs representing tertiary structures of enzymes from the BRENDA database; (4) Stochastic Block Model (SBM), a larger-scale dataset discussed separately in Sec. F.2 due to its distinct evaluation protocol.
We follow the standard experimental and evaluation settings from Jo et al. (2022), including the same train/test splits and the use of the Gaussian Earth Mover’s Distance (EMD) kernel for MMD computation, to ensure fair comparisons with baseline models — except for the Stochastic Block Model (SBM) dataset. We use node degree and spectral features of the graph Laplacian decomposition as hand-crafted input features. Baseline results are sourced from Jo et al. (2022); Kong et al. (2023) or reproduced using the corresponding publicly available code.
@@ E.3 Generic Graph Generation
Tab. 5 summarizes the key characteristics of the datasets utilized in this study. The table outlines the type of dataset, the total number of graphs, and the range of graph sizes ($|V|$). Additionally, it also provides the number of distinct node types and edge types for each dataset. Notably, the Community-small and Ego-small datasets contain relatively small graphs, whereas Enzymes, SBM and the molecular datasets (QM9 and ZINC250k) exhibit greater diversity in terms of graph size and complexity.
@@
Dataset Graph type #Graphs #Nodes
Node
types
Edge
types
@@
Appendix F Additional Experimental Results
F.1 Analysis of Diffusion Domain Choice
To investigate the impact of diffusion domain choice, we perform ablation experiments comparing two variants of HOG-Diff: one that operates directly in the adjacency matrix domain, and another in the Laplacian spectral domain. Both variants are trained under the same hyperparameter search space for a fair comparison.
As shown in Tab. 6, the spectral variant achieves comparable performance to the adjacency-based approach across most evaluation metrics on QM9 and ZINC250k. Despite similar results, we adopt the Laplacian spectral domain as the default diffusion space in HOG-Diff due to its theoretical and practical advantages. Specifically, the spectral domain offers greater efficiency and aligns naturally with core graph principles such as permutation invariance and signal concentration on low-dimensional manifolds.
Table 6: Comparison of diffusion domains in HOG-Diff.
[…]
QM9 Adjacency matrix 0.0004 0.264 99.08 100.00 95.90 67.78
Laplacian spectrum 0.0003 0.172 98.74 100.00 97.10 75.12
ZINC250k Adjacency matrix 0.006 4.259 96.75 100.00 99.78 99.98
Laplacian spectrum 0.001 1.633 98.56 100.00 99.96 99.53
F.2 Scalability Evaluation on Large Graphs
To further evaluate the scalability and robustness of HOG-Diff, we report results on the Stochastic Block Model (SBM) dataset in Tab. 7, which comprises graphs of larger scale and is evaluated under a distinct experimental protocol from that used in Sec. 4.2. The dataset consists of 200 synthetic graphs generated using the stochastic block model. The number of communities is uniformly sampled between 2 and 5, and the number of nodes within each community is uniformly sampled between 20 and 40. Edges are created with probabilities of 0.3 for intra-community connections and 0.05 for inter-community connections.
We adopt the evaluation setting introduced by Martinkus et al. (2022), including the same data splits, feature initialization, and the use of Total Variation (TV) distance to compute the MMDs. The TV kernel is adopted since it offers higher computational efficiency compared to the Earth Mover’s Distance (EMD) kernel, especially for large graphs. The baselines are sourced from (Vignac et al., 2023; Jo et al., 2024) or reproduced using the corresponding publicly available code. As shown in Tab. 7, HOG-Diff achieves competitive performance compared to the state-of-the-art.
Table 7: Generation results on the SBM dataset. The best and second-best results are highlighted in bold and underlined, respectively.
Method Deg. $\downarrow$ Clus. $\downarrow$ Orb. $\downarrow$ Spec. $\downarrow$ Avg. $\downarrow$
GraphRNN (You et al., 2018) 0.0055 0.0584 0.0785 0.0065 0.0372
GRAN (Liao et al., 2019) 0.0113 0.0553 0.0540 0.0054 0.0315
EDP-GNN (Niu et al., 2020) 0.0011 0.0552 0.0520 0.0070 0.0288
SPECTRE (Martinkus et al., 2022) 0.0015 0.0521 0.0412 0.0056 0.0251
GDSS (Jo et al., 2022) 0.0212 0.0646 0.0894 0.0128 0.0470
DiGress (Vignac et al., 2023) 0.0013 0.0498 0.0434 0.0400 0.0336
GruM (Jo et al., 2024) 0.0015 0.0589 0.0450 0.0077 0.0283
DeFoG (Qin et al., 2025) 0.0006 0.0517 0.0556 0.0054 0.0283
HOG-Diff (Ours) 0.0028 0.0500 0.0428 0.0043 0.0249
F.3 Rationale for Selecting Topological Filters
The filtering strategy in HOG-Diff is primarily driven by data statistics and complemented by domain knowledge to ensure meaningful choices.
In molecular generation experiments, we employ 2-cells as guides because they are ubiquitous in molecular graphs and capture critical chemical information, e.g., functional groups. In particular, 2-cells play a crucial role in determining the three-dimensional conformation, electron distribution, and target binding mode of compounds, making them highly informative. Consequently, Tab. 1 shows that conditioning on 2-cells alone provides strong guidance and achieves competitive results across all metrics.
In contrast, cells with dimension >2>2 are extremely sparse in the standard benchmarks (see Tab. 8). Identifying them requires increasingly complex preprocessing, such as detecting candidate higher-order topological structures and verifying their validity, which introduces non-trivial computational overhead. Moreover, given that these benchmarks contain only a few hundred nodes, guides with higher-dimensional cells would cover less than 0.05% of possible structures, introducing both unnecessary computation and a risk of overfitting without evident benefit. In practice, 2-cells already capture virtually all of the higher-order structure present in these datasets.
We further tested guiding molecular generation with 2-simplices but observed worse performance. This is expected, as summarized in Tab. 8, higher-dimensional simplices are extremely rare in molecular datasets (i.e., QM9 and ZINC250k), while 2-cells are comparatively more abundant. Moreover, most functional structures in molecules do not satisfy the requirements of simplices. These findings confirm that aligning the guide with the natural higher-order structures of the data improves both training efficiency and sample quality, whereas mismatched guides provide limited benefit.
In contrast, generic graphs (particularly those in social or biological domains) often exhibit diverse simplicial structures, making simplicial filtering more suitable. This demonstrates the adaptability of HOG-Diff across domains.
Overall, the proposed framework is most effective when the intermediate skeletons reflect structures naturally present in the data. This aligns with prior work in graph representation learning, which shows that the benefits of higher-order representations are most pronounced in datasets rich in such structures (Huang et al., 2024). Indeed, many real-world networks—including social, biological, and citation networks—exhibit abundant higher-order cells/simplices. Thus, many real-world tasks naturally fall into the “topologically rich” regime where our method is particularly advantageous.
Table 8: Average counts of higher-order structures across datasets. “–” indicates computation exceeded 10h.
Dataset 2-simplices 3-simplices 4-simplices 2-cells 3-cells
Community-small 27.45 8.24 0.76 40.92 0.00
Ego-small 4.61 1.64 0.35 5.66 0.00
Enzymes 25.93 3.05 0.01
SBM 441.25 90.88 4.42
QM9 0.47 0.00 0.00 1.84 0.10
ZINC250k 0.06 0.00 0.00 2.77 0.00
F.4 De Novo Molecule Generation
Figure 8: Visualization of chemical space distributions. Molecular representations are obtained using Morgan fingerprints and subsequently visualized through dimensionality reduction with Uniform Manifold Approximation and Projection (UMAP). Contour lines denote the probability distributions of reference (ref) and generated (gen) molecules. (a) illustrates the distributional shift across QM9 (green) and ZINC250k (blue), while (b) and (c) show that our generative model faithfully captures the distinct distributions of each dataset.
To visually assess the capability of the proposed method in molecular generation, Fig. 8 compares distributional results on the molecular datasets. Specifically, we first calculate the Morgan fingerprints (Rogers & Hahn, 2010) of all molecules, which have been widely utilized in drug discovery for capturing structural information. Subsequently, we apply Uniform Manifold Approximation and Projection (UMAP), a nonlinear dimensionality-reduction method that preserves local similarities, to embed the fingerprints into two dimensions and plot the resulting distributions.
On the QM9 dataset, the distributions generated by our model closely align with those of the reference set. On the more complex ZINC250k dataset, the generated distributions show slight deviations but remain well aligned. These results provide an intuitive demonstration of strong generative performance and complement the findings in Tab. 1.
We provide the standard deviation results and the additional validity (Val.) metric in Tabs. 9 and 10. Baseline results are sourced from Jo et al. (2022); Kong et al. (2023) or reproduced using the corresponding publicly available code.
Table 9: Comparison of different methods on QM9. We report the means and standard deviations of 3 runs. Asterisks (*) indicate that the source did not report standard deviations. The best results for the first three metrics are highlighted in bold.
[…]
Val. w/o
[…]
GraphArm* 0.002 1.220 90.25 100.00 95.62 70.39
MiCaM* 0.001 1.045 99.93 100.00 93.89 83.25
[…]
SPECTRE* 0.163 47.960 87.30 100.00 35.70 97.28
GSDM* 0.003 2.650 99.90 100.00 - -
[…]
DiGress* 0.0005 0.360 99.00 100.00 96.66 33.40
[…]
CatFlow* - 0.441 99.81 100.00 99.95 -
[…]
Table 10: Comparison of different methods on ZINC250k. We report the means and standard deviations of 3 runs. Asterisks (*) indicate that the source did not report standard deviations. The best results for the first three metrics are highlighted in bold.
[…]
Val. w/o
[…]
GraphArm* 0.055 16.260 88.23 100.00 99.46 100.00
MiCaM* 0.166 31.495 100.00 100.00 88.48 99.98
[…]
SPECTRE* 0.109 18.440 90.20 100.00 67.05 100.00
GSDM* 0.017 12.956 92.70 100.00 - -
[…]
DiGress* 0.082 23.060 91.02 100.00 81.23 100.00
[…]
CatFlow* - 13.211 99.95 99.99 100.00 -
[…]
F.5 Visualization Results